Status Report

Hubble Space Telescope Daily Report #2828 14 Mar 2001

By SpaceRef Editor
March 14, 2001
Filed under ,

HUBBLE SPACE TELESCOPE

DAILY REPORT #2828

PERIOD COVERED: 0000Z (UTC) 03/13/01 – 0000Z (UTC) 03/14/01

Daily Status Report as of 073/0000Z

1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED:

1.1 Completed Three Sets of WF/PC-2 8817 (Cycle 9 INTFLAT Sweeps and
Linearity Test)

The WF/PC-2 was used to obtain INTFLAT observations in order to
monitor the pixel to pixel flatfield response and provide a linearity
check. The INTFLAT sequences, to be done once during the year, are similar
to those from the Cycle 8 program 8448. The images will provide a backup
database in the event of complete failure of the VISFLAT lamp as well as
allow monitoring of the gain ratios. The sweep is a complete set of
internal flats, cycling through both shutter blades and both gains. The
linearity test consists of a series of INTFLATs in F555W, in each gain and
each shutter. There was no anomalous activity.

1.2 Completed FGS/1 8731 (A High Angular Resolution Survey of the Most
Massive Stars in the SMC)

Fine Guidance Sensor #1 was used to conduct high angular resolution
observations of some of the SMC’s hottest and most luminous objects by
conducting a comprehensive survey of nearly every spectral type in the
upper portion of the HR diagram. Binary or multiple star systems will be
detected down to an unprecedented level, more than one order of magnitude
better than possible with WFPC2. The targets selected include a
representative list of normal Main Sequence O-stars and their evolved
descendents, namely supergiants, hypergiants, LBVs, and WRs, many of which
have been previously observed by HST’s spectrometers for purposes of
cosmological calibrations. The observations completed, and no problems
were reported.

1.3 Completed STIS/CCD/MA1 8889 (Tortured Coronae in the Rapid Breaking
Zone)

The Space Telescope Imaging Spectrograph (CCD and MA1) was used to
perform observations in conjunction with deep Chandra X-ray Observatory
high energy transmission grating {HETGS} exposures of three key G0-G5
giants that will address fundamental questions concerning their hot
magnetized coronae by determining temperatures and densities, probing
chemical fractionation processes, and exploring plasma dynamics. The
observations completed nominally.

1.4 Completed WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive
Proposal Continuation)

The WF/PC-2 was used to perform a generic target version of the
WFPC2 Archival Pure Parallel program. The program will be used to take
parallel images of random areas of the sky, following the recommendations
of the Parallels Working Group. The proposal completed with no reported
problems.

1.5 Completed STIS/MA1 8862 (FUV-MAMA Cycle 9 Flats)

The Space Telescope Imaging Spectrograph (MA1) was used to obtain
FUV-MAMA observations of the internal Krypton lamp to construct an FUV flat
applicable to all FUV modes. No problems were encountered.

1.6 Completed WF/PC-2 8816 (Cycle 9 UV Earthflats)

The WF/PC-2 was used to obtain sequences of Earth streak flats to
improve the quality of pipeline flat fields for the WFPC2 UV filter set and
in order to monitor flat field stability. There were no reported problems.

1.7 Completed WF/PC-2 8686 (LINERs in Early-type Galaxies: Ionized by
the UV-upturn Population?)

The WF/PC-2) was used to investigate the presence of extended
ionized gas with LINER spectra in a large fraction {60-70%} of early-type
galaxies. The observations completed with no reported anomalies.

1.8 Completed STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic
Shear Observations)

The Space Telescope Imaging Spectrograph (CCD) was used to probe
the distortion of light bundles from distant galaxies, looking at the
statistical properties of the intervening inhomogeneous {dark} matter
distribution. The proposal completed nominally.

2.0 FLIGHT OPERATIONS SUMMARY:

2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 11

Successful: 11

Scheduled Re-acquisitions: 4

Successful: 4

2.2 FHST Updates:

Scheduled: 21

Successful: 21

2.3 Operations Notes:

Operations personnel, in addition to supporting on-going operations
from the Science Institute, participated in GSFC Internal Simulation #6
(EVA-5) from 8 a.m. until approximately 5:30 p.m.

The SSR EDAC error counter was cleared twice per ROP SR-1A.

The STIS MCE-1 reset at 072/155404Z while the low voltage was on
and while outside any SAA interval. The STIS flight software error counter
was cleared at 072/1608Z, using ROP NS-12. MAMA-1 was recovered at
072/1620Z by normal SMS commanding when the high voltage was turned on.

3.0 SIGNIFICANT FORTHCOMING EVENTS:

Continuation of normal science observations and calibrations.

SpaceRef staff editor.